OpenStudy (anonymous):

(x+0,2)^2-(x-0,2)^2=?

OpenStudy (turingtest):

are you squaring ordered pairs?

OpenStudy (anonymous):

i got 0

OpenStudy (anonymous):

The book says the answer is 0.8x but I cant seem to make it work on paper

OpenStudy (turingtest):

(x+0.2)^2-(x-.02)^2 =(x^2+0.4x+0.04)-(x^2-0.4x+0.04) =x^2+0.4x+0.04-x^2+0.4x-0.04 =0.8x There is no such thing as multiplying ordered pairs, I was just making fun of the fact that you put commas instead of points in your question.